# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

### v2

@@ 5.3 Proxies & Disagreement Correlation
For Normative Uncertainty, we choose KeyCase as a proxy following Xu et al. 2023b. The court annually chooses a set of “key cases”, which often deal with complex and novel legal issues. Given the absence of established legal standards for them, they often generate controversy and disagreement among experts. We hypothesize that judge tend to disagree more in “key cases”.
Specific Legal Context via two proxy variables:
HighRep (High Reputation Country): Previous legal and political science work indicates ECtHR Judges tend to have more split votes in cases where the defending country is regarded as High Reputation Country [^10] Countries with higher democracy index. See App K. Dothan (2014) .
VulnApplicant (Vulnerability of Applicant): the ECtHR court adapts convention standards to meet vulnerable individuals’ needs despite vulnerability not being defined in the Convention. It remains underdefined within the context of ECtHR because the judges do not explain the process though which they identify an individual as vulnerable Heri (2021), which can lead to divergent legal opinions. We retrieve relevant information from the VECHR dataset Xu et al. (2023a) and categorize a case as 1 if its applicant is regarded as vulnerable. [^11] Only annotation of cases under Article 3 are available in VECHR before 2016. Hence our t-test of VulnApplicant is only on cases of Article 3. App L offers more explanation on Vulnerability in ECtHR jurisprudence.
Subjectivity refers to the disagreement due to personal opinions and values. We propose that the 17-judge Grand Chamber is more prone to disagreement than the 7-judge chamber. This is due to the inherent pluralism of human values in larger groups, which naturally fosters a broader range of viewpoints and thus, increased subjectivity. As explained in sec. 5.1.2, modeling specific ideological dimensions of individual judges is out of scope for this work. We account for some limited political dynamics by including the democratic score of the state against which the claim is brought (the HighRep variable above).
@@ 5.3.1 Do proxy measures correlate with judges’ votes?
Tab 1 shows significantly lower agreement among judges’ votes for HighRep and VulnApplicant (p-value < 0.05). In other words, specific contexts of case facts correlate more with disagreement than Normative Uncertainty. Judges express disagreement related to how facts fit the norm because of circumstances related to the social standing of the applicant and the political situation of the defending State. This is an inherent characteristic of adjudication where judges qualify the context so as to fit the norm. The inclusion of a category related to the legal norm’s scope could have provided further insights.
@@
Interestingly, we found less disagreement among judges on KeyCase. Experts suggest this may stem from the Court’s assigning the importance score ex post, after interpretation and agreement are reached. The Court is more likely to designate a consensus case as KeyCase to maintain coherent jurisprudence, rather than a controversial case that could invite future applicants to base cases on disputed facts and dissenting opinions. As reflected in our DoF framework (App A), Case to Law uncertainty inherently perpetuates Law to Case uncertainty.
@@ 6.1 Experimental Setup of COC
We assess the alignment of a model’s prediction difficulty [^12] Another approach would be assessing a model’s prediction uncertainty. Refer to App O for an explanation of our decision to abstain from using Uncertainty evaluation methods with judge disagreement, specifically COC models trained on ECtHR Task A|B. To address input length constraints, we employ a LegalBERT Chalkidis et al. (2020) variant of the hierarchical attention model [^13] See App P for model structure and implementation. Yang et al. (2016), as adapted from Santosh et al. 2022.
Task Benchmark task ECtHR A|B (Violation Identification given Allegation) identifies the set of violated convention article(s) from the textual case facts and list of allegedly violated articles.
Models Santosh et al. 2022 encodes allegation information as a multi-hot vector concatenated with the text representation. By contrast, in this study, we textify allegation information as “Alleged Article: X”, where is X is a comma-separated list of articles and prepend to the input text. We employ the LegalBERT nlpaueb/legal-bert-base-uncased as the backbone in hierarchical models [^14] To explore the impact of domain-specific pre-training, we did also experiment on BERT. See App M..
Metrics Our evaluation of COC performance employs the micro-F1 (mic-F1), macro-F1 (mac-F1), and hard-macro-F1 (hm-F1; Santosh et al. 2022). The hm-F1 is calculated as the mean F1-score for each article, where cases with that article having been violated are considered as positive instances, and cases with that article being alleged but not found to be violated as negative instances, resulting in a smaller pool of more difficult negatives.
COC Performance We run experiments across three random seeds. The hierarchical LegalBERT achieves 73.55 ± 0.62 for mic-F1, 68.75 ± 1.13 for mac-F1 and 68.03 ± 1.29 for hm-F1. Although the mic-F1 and mac-F1 performances closely align with those reported in Santosh et al. (2022); Xu et al. (2023b), differences of $\sim$ 2%, the hm-F1 performance demonstrate a noteworthy improvement of $\sim$10%. [^15] This can be partly attributed to our correction of labels (see § 4). Some variation may also be due to randomness in weight initialization. See our discussion in App N Importantly, hm-F1 exhibits particular sensitivity to allegation information. For the rest of this paper, we adopt hm-F1 as our primary metric for COC performance.
@@ 6.2 Difficulty of SV Cases in COC
Experiment Settings: We use the COC fine-tuned models from Sec 6.1 as the input-target model $g$. We fine-tune an architecturally identical model $g^{\prime}$ with the input replaced by $\emptyset$. Subsequently, we partition the test set into two distinct partitions of SV and unanimous cases, respectively. Our expectation is that the SV subset will exhibit lower average PVI compared to the unanimous subset. We use entropy of the judges’ vote distribution as an estimator of case difficulty: the more the judges disagree with each other, the harder a case is.
Results: Tab 2 presents the average COC performance (hm-F1) and model-perceived difficulty (mean PVI scores). The model exhibits lower hm-F1 and PVI scores for SV cases compared to unanimous cases. An independent t-test on the average PVI scores between the two groups reveals a significant difference ($p<0.05$). This indicates that SV cases are more challenging for the models than unanimous cases, aligning with human perceptions of difficulty. We also calculate the Pearson correlation coefficients between PVI scores and the entropy of judges’ vote distribution of SV cases. LegalBERT has a correlation coefficient (r-value) of -0.068 and a p-value of 0.48. The results reveal a negative correlation between PVI scores and the degree of disagreement among judges, consistent with our expectations. However, the observed correlation is very weak ($|r|<0.1$) and statistically insignificant ($p>0.05$). This suggests that, while the model can capture differences in difficulty between unanimous and SV cases, it struggle to accurately represent the nuanced degree of judges’ disagreement. Therefore, we next evaluate the models’ confidence and soft-label training.
@@ 6.3 Calibration of SV Cases
hm-F1 $\uparrow$ ECE $\downarrow$ DistCE $\downarrow$ count
/ TS soft / TS soft / TS soft
[…]
Table 3: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std ($\pm$) on test set. "/": COC finetuned models in § 6.1; "soft": models fine-tuned with soft-loss; "TS": models after Temperature Scaling. Results shown over 3 random seeds. See Tab 8 in App S for performance on dev set.
@@ 6.3.1 Evaluation Metrics
Figure 3: The distributions over probabilities for class 1 of the models vs human vote distributions (row 1) and distCE (row 2). See Fig 8 in App S for more figures comparing human uncertainty to model uncertainty
@@ 6.3.2 Calibration Methods
For confidence calibration, we employ Temperature Scaling (TS; Guo et al. 2017). This simple yet widely used post-hoc method uses a single temperature parameter $t$ to scale the output logits of a classifier. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. For human calibration, we adopt the approach from Peterson et al. 2019; Uma et al. 2020 by finetuning with a Soft Loss Function. [^16] See App Q for details of the Soft Loss Function. During training, models are exposed to “soft labels” derived from the judges’ vote distributions, serving as target distributions in a cross-entropy loss function.
Expected Calibration Error (ECE) is the most often used metric for confidence calibration Naeini et al. (2015); Guo et al. (2017). A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true probabilities. Refer to the App R for further details on ECE.
DistCE by Baan et al. 2022 measures human calibration, which can be calculated as $\mathrm{DistCE}(x)=1/2\|\mathbf{q}-\mathbf{p}\|$, where $q$ is the vote distribution and $p$ is the model’s predictive distribution.
@@ 6.3.3 Results & Discussion
Tab 3 confirms that SV cases are indeed the most challenging, reflected by the lowest COC performance (hm-F1). They also exhibit a higher degree of miscalibration (higher ECE) and misalignment with human responses (higher distCE). Applying TS greatly reduces ECE, without negatively impacting hm-F1, enhancing confidence calibration.
However, the models remain misaligned with the human vote distribution, as the high DistCE scores show and Fig 3 illustrates. Soft-loss tuning slightly improves the alignment between the model’s prediction confidence and the human vote distribution (lower DistCE). Soft LegalBERT exhibits noticeably fewer instances of over-confident predictions when judges do not unanimously agree, as depicted on near-0/1 probability portions of Fig 3b, in contrast to the original model shown in Fig 3a. TS models as expected reduce overly miscalibrated instances (drop in tail) but they exhibit substantially lower DistCE scores on split-vote cases than soft models. This may be attributed to an overly aggressive temperature, resulting in a more uniform output distribution (see Fig 3c, with a temperature of 5.5). Moreover, judge votes often exhibit a quasi-bimodal distribution, with many split votes caused by a single judge dissenting pro/con the finding of a violation (see discussion on single dissenting votes in § 4). Further, the DistCE score histograms (Fig 3d-f) illustrate that soft-tuning (Fig 3e) improves human alignment by reducing instances of extreme miscalibrations in the right tail, as compared to the original model in Fig 3d. It is noteworthy that the TS mode (Fig 3f) shifts the distribution further to the left by reducing the right tail. However, this is at the expense of predictions that perfectly align with human judgement probabilities, as evidenced by fewer instances of DistCE scores below 0.15 in the TS model compared to the Soft model. Therefore, despite TS models potentially displaying lower DistCE scores than soft models, they do not provide an optimal fit. It is crucial to analyze the distributions prior to drawing conclusions, especially as we have shown in cases with split votes, where distributions are bimodal.
@@ 7 Conclusion
Appendix A DoF of Sources of Disagreement in Legal Decision Process
To better capture the circular nature of legal decision process, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the recurrent nature of this process and the continuous reproduction of the main normative model over time.
Figure 5: Screenshot of Judges’ vote information in HUDOC
Appendix B Information about Judge’ Votes in HUDOC
@@
Appendix C Quality Validation of SV-ECHR
@@
Appendix D Dataset Statistics
@@
Appendix E Correction of Inconsistent Metadata
@@
Tab 5 shows the statistics about our correction of meta information. Notably, apart from the inconsistent allegation information pointed out by Xu et al., 2023b. We further found approximately 2% inconsistent violation information in the train set during our quality control process. Most of the cases with so-called striked-out cases which are difficult to parse. We also find finetuning on the corrected train set improves model’s COC performance (§ 6.1) as discussed in App N.
Appendix F Entropy of Judges’ Vote Distribution
Fig 6 display the histogram of the entropy of each pair’s vote distribution. The entropy is calculated with $\mathbf{H}(\mathbf{p})=-\sum_{i\in\mathcal{C}}p_{i}\log\left(p_{i}\right)$ and $p_{i}=\frac{n_{i}}{\sum_{j\in\mathcal{C}}n_{j}}$ , where C is the label category set [violation, non-violation] and $n_{i}$ is the number of judge voting for category i.
Appendix G Debates in Legal Scholarship about Single Judge Dissenting Vote
Tab 4 reveals that $\sim$ 60% of chamber votes (involving 7 judges) exhibit entropy of around 0.4, indicating that only one judge vote differently from the remaining 6 judges (entropy([1,6]) $\approx$ 0.4). A similar pattern is observed for Grand Chamber cases involving 17 judges. This Single Judge Dissenting pattern aligns with ongoing debates in legal scholarship. For instance, Fobbe 2022 examined the number of dissenting opinions in decisions from the International Court of Justice (ICJ). Their results indicate a significant proportion of unanimous decisions, followed by a monotonously decreasing number of dissents. Some legal scholars support the view that diverging opinions do not signal a division of the bench inside the Court over the scope of protection of rights, even less a lowering of standards of protection. While other scholars, support the contrary view according to which dissenting opinions send a signal of walking-back in terms of effective protection. Tab 4 and Fobbe 2022’s finding actually shows that the case law does not consistently but rather seldomly give rise to dissenting opinions. Moreover, one can not ignore the National Judge’s dissenting vote (App I). It is difficult to assess what is the weight of a single judge dissenting on how the court will decide cases. Often times, dissenting opinions are not followed upon. What is of relevance is to determine a yardstick against which to measure disagreement in human decisions and this is what we are doing.

### v3

@@ 5.3 Proxies & Disagreement Correlation
For Normative Uncertainty, we choose KeyCase as a proxy following Xu et al. 2023b. The court annually chooses a set of “key cases”, which often deal with complex and novel legal issues. Given the absence of established legal standards for them, they often generate controversy and disagreement among experts. We hypothesize that judges tend to disagree more in “key cases”.
Specific Legal Context
via two proxy variables:
HighRepCountry (High Reputation Country): Previous legal and political science work indicates ECtHR judges tend to have more split votes in cases where the defending country is regarded as a High Reputation Country***Countries with higher democracy index. See App J. Dothan (2014) .
VulnApplicant (Vulnerability of Applicant): the ECtHR court adapts convention standards to meet vulnerable individuals’ needs despite vulnerability not being defined in the Convention. It remains underdefined within the context of ECtHR because the judges do not explain the process though which they identify an individual as vulnerable Heri (2021), which can lead to divergent legal opinions. We retrieve relevant information from the VECHR dataset Xu et al. (2023a) and categorize a case as 1 if its applicant is regarded as vulnerable.***Only annotation of cases under Article 3 are available in VECHR before 2016. Hence our t-test of VulnApplicant is only on cases of Article 3. App K offers more explanation on Vulnerability in ECtHR jurisprudence.
Subjectivity
refers to the disagreement due to personal opinions and values. We propose that the 17-judge Grand Chamber is more prone to disagreement than the 7-judge Chamber. This is due to the inherent pluralism of human values in larger groups, which naturally fosters a broader range of viewpoints and thus, increased subjectivity. As explained in sec. 5.1.2, modeling specific ideological dimensions of individual judges is out of scope for this work. We account for some limited political dynamics by including the democratic score of the state against which the claim is brought (the HighRep variable above).
@@ 5.3.1 Do proxy measures correlate with judges’ votes?
Tab 1 shows significantly lower agreement among judges’ votes for HighRepCountry and VulnApplicant (p-value < 0.05). In other words, specific contexts of case facts correlate more with disagreement than Normative Uncertainty. Judges express disagreement related to how facts fit the norm because of circumstances related to the social standing of the applicant and the political situation of the defending State. This is an inherent characteristic of adjudication where judges qualify the context so as to fit the norm. The inclusion of a category related to the legal norm’s scope could have provided further insights.
@@
Interestingly, we found less disagreement among judges on KeyCase. Experts suggest this may stem from the Court’s assigning the importance score ex post, after interpretation and agreement are reached. The Court is more likely to designate a consensus case as KeyCase to maintain coherent jurisprudence, rather than a controversial case that could invite future applicants to base cases on disputed facts and dissenting opinions. As reflected in our DoF framework (LABEL:app:dof), Case-to-Law uncertainty inherently perpetuates Law-to-Case uncertainty.
hm-F1 $\uparrow$ ECE $\downarrow$ DistCE $\downarrow$ count
/ TS soft / TS soft / TS soft
[…]
Table 2: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std ($\pm$) on test set. "/": COC finetuned models in § 6.1; "soft": models fine-tuned with soft-loss; "TS": models after Temperature Scaling. Results shown over 3 random seeds. See Tab 8 in App Q for performance on dev set.
Figure 4: The distributions over probabilities for class 1 of the models vs human vote distributions (row 1) and distCE (row 2). See Fig 8 in App Q for more figures comparing human uncertainty to model uncertainty
@@ 6.1 Experimental Setup of COC
We assess the alignment of a model’s prediction difficulty***Another approach would be assessing a model’s prediction uncertainty. Refer to § 2 for an explanation of our decision to abstain from using Uncertainty evaluation methods with judge disagreement, specifically COC models trained on ECtHR Task A|B. To address input length constraints, we employ a LegalBERT Chalkidis et al. (2020) variant of the hierarchical attention model***See App N for model structure and implementation. Yang et al. (2016), as adapted from Santosh et al. 2022.
Task
Benchmark task ECtHR A|B (Violation Identification given Allegation) identifies the set of violated convention article(s) from the textual case facts and list of allegedly violated articles.
Models
Santosh et al. 2022 encodes allegation information as a multi-hot vector concatenated with the text representation. By contrast, in this study, we textify allegation information as “Alleged Article: X”, where is X is a comma-separated list of articles and prepend to the input text. We employ the LegalBERT (nlpaueb/legal-bert-base-uncased; Chalkidis et al. 2020) as the backbone in hierarchical models.***To explore the impact of domain-specific pre-training, we did also experiment on BERT. See App L.
Metrics
Our evaluation of COC performance employs the micro-F1 (mic-F1), macro-F1 (mac-F1), and hard-macro-F1 (hm-F1; Santosh et al. 2022). The hm-F1 is calculated as the mean F1-score for each article, where cases with that article having been violated are considered as positive instances, and cases with that article being alleged but not found to be violated as negative instances, resulting in a smaller pool of more difficult negatives.
COC Performance
We run experiments across three random seeds. The hierarchical LegalBERT achieves 73.55 ± 0.62 for mic-F1, 68.75 ± 1.13 for mac-F1 and 68.03 ± 1.29 for hm-F1. Although the mic-F1 and mac-F1 scores closely align with those reported in Santosh et al. (2022); Xu et al. (2023b), differences of $\sim$ 2%, the hm-F1 performance demonstrate a noteworthy improvement of $\sim$10%.***This can be partly attributed to our correction of labels (see § 4). Some variation may also be due to randomness in weight initialization. See our discussion in App M Importantly, hm-F1 exhibits particular sensitivity to allegation information. For the rest of this paper, we adopt hm-F1 as our primary metric for COC performance.
@@ 6.2 Difficulty of SV Cases in COC
Experiment Settings
We use the COC fine-tuned models from Sec 6.1 as the input-target model $g$. We fine-tune an architecturally identical model $g^{\prime}$ with the input replaced by $\emptyset$. Subsequently, we partition the test set into two distinct partitions of SV and unanimous cases, respectively. Our expectation is that the SV subset will exhibit lower average PVI compared to the unanimous subset. We use entropy of the judges’ vote distribution as an estimator of case difficulty: the more the judges disagree with each other, the harder a case is.
Results
Tab 3 presents the average COC performance (hm-F1) and model-perceived difficulty (mean PVI scores). The model exhibits lower hm-F1 and PVI scores for SV cases compared to unanimous cases. An independent t-test on the average PVI scores between the two groups reveals a significant difference ($p<0.05$). This indicates that SV cases are more challenging for the models than unanimous cases, aligning with human perceptions of difficulty. We also calculate the Pearson correlation coefficients between PVI scores and the entropy of judges’ vote distribution of SV cases. LegalBERT has a correlation coefficient (r-value) of -0.068 and a p-value of 0.48. The results reveal a negative correlation between PVI scores and the degree of disagreement among judges, consistent with our expectations. However, the observed correlation is very weak ($|r|<0.1$) and statistically insignificant ($p>0.05$). This suggests that, while the model can capture differences in difficulty between unanimous and SV cases, it struggle to accurately represent the nuanced degree of judges’ disagreement. Therefore, we next evaluate the models’ confidence and soft-label training.
@@ 6.3.2 Calibration Methods
For confidence calibration, we employ Temperature Scaling (TS; Guo et al. 2017). This simple yet widely used post-hoc method uses a single temperature parameter $t$ to scale the output logits of a classifier. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. For human calibration, we adopt the approach from Peterson et al. 2019; Uma et al. 2020 by finetuning with a Soft Loss Function.***See App O for details of the Soft Loss Function. During training, models are exposed to “soft labels” derived from the judges’ vote distributions, serving as target distributions in a cross-entropy loss function.
Expected Calibration Error
(ECE) is the most often used metric for confidence calibration Naeini et al. (2015); Guo et al. (2017). A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true probabilities. Refer to the App P for further details on ECE.
DistCE
by Baan et al. 2022 measures human calibration, which can be calculated as $\mathrm{DistCE}(x)=1/2\|\mathbf{q}-\mathbf{p}\|$, where $q$ is the vote distribution and $p$ is the model’s predictive distribution.
@@ 6.3.3 Results & Discussion
Tab 2 presents the experimental results. We make the following observations:
i) SV cases are indeed the most challenging.
It is reflected by their lowest COC performance (hm-F1). They also exhibit a higher degree of miscalibration (higher ECE) and misalignment with human responses (higher distCE).
ii) Applying TS greatly enhanced confidence calibration.
Tab 2 presents that applying TS has greatly reduced the ECE score, without negatively impacting COC performance (hm-F1).
iii) The models remain misaligned with the human vote distribution.
Tab 2 shows that soft-loss tuning only slightly improves the alignment between the model’s prediction confidence and the human vote distribution (lower DistCE). Fig 4 illustrates that Soft LegalBERT exhibits noticeably fewer instances of over-confident predictions when judges do not unanimously agree, as depicted on near-0/1 probability portions of Fig 4b, in contrast to the original model shown in Fig 4a. On split-vote cases, TS models exhibit substantially lower DistCE scores than soft models. However, we do not consider that TS provides a better human calibration. The lower DistCE scores may be attributed to an overly aggressive temperature, resulting in a more uniform output distribution (see Fig 4c, with a temperature of 5.5). Moreover, judge votes often exhibit a quasi-bimodal distribution, with many split votes caused by a single judge dissenting pro/con the finding of a violation (see discussion on single dissenting votes in § 4). Further, the DistCE score histograms (Fig 4d-f) illustrate that soft-tuning (Fig 4e) improves human alignment by reducing instances of extreme miscalibrations in the right tail, as compared to the original model in Fig 4d. It is noteworthy that the TS model (Fig 4f) shifts the distribution further to the left by reducing the right tail. However, this is at the expense of predictions that perfectly align with human judgement probabilities, as evidenced by fewer instances of DistCE scores below 0.15 in the TS model compared to the Soft model. Therefore, despite TS models potentially displaying lower DistCE scores than soft models, they do not provide an optimal fit. It is crucial to analyze the distributions prior to drawing conclusions, especially as we have shown in cases with split votes, where distributions are bimodal.
@@ 7 Conclusion
Acknowledgements
We would like to thank the anonymous reviewers for their valuable comments and suggestions. We would also like to thank the members of the MaiNLP/CIS research lab for their thoughtful feedback. We acknowledge the support of our funders. BP is supported by the ERC Consolidator Grant No. 101043235.
Appendix A Information about Judge’ Votes in HUDOC
@@
Appendix B Quality Validation of SV-ECHR
@@
Appendix C Dataset Statistics
@@
Appendix D Correction of Inconsistent Metadata
@@
Tab 5 shows the statistics about our correction of case metadata. Notably, apart from the inconsistent allegation information pointed out by Xu et al., 2023b. We further found approximately 2% inconsistent violation information in the train set during our quality control process. Most of the cases are so-called ‘striked-out’ cases which are difficult to parse. We also find finetuning on the corrected train set improves model’s COC performance (§ 6.1) as discussed in App M.
Appendix E Entropy of Judges’ Vote Distribution
Fig 6 display the histogram of the entropy of each pair’s vote distribution. The entropy is calculated as $\mathbf{H}(\mathbf{p})=-\sum_{i\in\mathcal{C}}p_{i}\log\left(p_{i}\right)$ and $p_{i}=\frac{n_{i}}{\sum_{j\in\mathcal{C}}n_{j}}$ , where $\mathcal{C}$ is the label category set [violation, non-violation] and $n_{i}$ is the number of judges voting for category $i$.
Appendix F Debates in Legal Scholarship about Single Judge Dissenting Vote
Tab 4 reveals that $\sim$ 60% of chamber votes (involving 7 judges) exhibit entropy of around 0.4, indicating that only one judge vote differently from the remaining 6 judges (entropy([1,6]) $\approx$ 0.4). A similar pattern is observed for Grand Chamber cases involving 17 judges. This Single Judge Dissenting pattern aligns with ongoing debates in legal scholarship. For instance, Fobbe 2022 examined the number of dissenting opinions in decisions from the International Court of Justice (ICJ). Their results indicate a significant proportion of unanimous decisions, followed by a monotonously decreasing number of dissents. Some legal scholars support the view that diverging opinions do not signal a division of the bench inside the Court over the scope of protection of rights, even less a lowering of standards of protection. Others support the contrary view according to which dissenting opinions send a signal of walking-back in terms of effective protection Helfer and Voeten (2020). Tab 4 and Fobbe 2022’s finding actually shows that the case law does not consistently, but rather seldomly, give rise to dissenting opinions. Moreover, one can not ignore the national judge’s dissenting vote (App H). It is difficult to assess the weight of a single judge dissenting on the decision of the court as a whole. Often times, dissenting opinions are not followed up on. Our focus here is limited to developing a yardstick against disagreement in human decisions can be measured.